1.      Find the range, the standard deviation, and the variance for the given samples. Round noninteger results to the nearest tenth.

       a. 1, 2, 5, 7, 8, 19, 22

       b. 3, 4, 7, 11, 12, 12, 15, 16

       c. 2.1, 3.0, 1.9, 1.5, 4.8

       d. 5.2, 11.7, 19.1, 3.7, 8.2, 16.3

       e. 48, 91, 87, 93, 59, 68, 92, 100, 81

1
Expert's answer
2022-05-03T03:52:49-0400

a.

We have sample size n=7.

The smallest data value is 1 and the largest is 22. The range is:


"22-1=21"

Mean = 

"\bar{x}=\dfrac{1+2+5+7+8+19+22}{7}=9.15"


Variance


"s^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n-1}=67.1"


Standard deviation

"s=\sqrt{s^2}=8.2"



b.

We have sample size n=8.

The smallest data value is 3 and the largest is 16. The range is:


"16-3=13"


Mean = 

"\bar{x}=\dfrac{3+4+7+11+12+12+15+16}{8}=10"


Variance


"s^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n-1}=23.4"


Standard deviation

"s=\sqrt{s^2}=4.8"



c.

We have sample size n=5.

The smallest data value is 1.5 and the largest is 4.8. The range is:


"4.8-1.5=3.3"


Mean = 

"\bar{x}=\dfrac{2.1+3.0+1.9+1.5+4.8}{5}=2.7"


Variance


"s^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n-1}=1.7"


Standard deviation

"s=\sqrt{s^2}=1.3"



d.

We have sample size n=6.

The smallest data value is 3.7 and the largest is 19.1. The range is:


"19.1-3.7=15.4"


Mean = 

"\bar{x}=\dfrac{5.2+11.7+19.1+3.7+8.2+16.3}{6}=10.7"


Variance


"s^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n-1}=37.7"


Standard deviation

"s=\sqrt{s^2}=6.1"



e.

We have sample size n=9.

The smallest data value is 48 and the largest is 100. The range is:


"100-48=52"


Mean = 

"\bar{x}=\dfrac{1}{9}(48+91+87+93+59"




"+68+92+100+81)=79.9"

Variance


"s^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n-1}=311.6"


Standard deviation

"s=\sqrt{s^2}=17.7"